# Enhancing Transformation from Natural Language to Signal Temporal Logic Using LLMs with Diverse External Knowledge

### v1

@@ Abstract
Temporal Logic (TL), especially Signal Temporal Logic (STL), enables precise formal specification, making it widely used in cyber-physical systems such as autonomous driving and robotics. Automatically transforming NL into STL is an attractive approach to overcome the limitations of manual transformation, which is time-consuming and error-prone. However, due to the lack of datasets, automatic transformation currently faces significant challenges and has not been fully explored. In this paper, we propose an NL-STL dataset named STL-Diversity-Enhanced (STL-DivEn), comprising 16,000 samples enriched with diverse patterns. To develop the dataset, we first manually create a small-scale seed set of NL-STL pairs. Next, representative examples are identified through clustering and used to guide large language models (LLMs) in generating additional NL-STL pairs. Finally, diversity and accuracy are ensured through rigorous rule-based filters and human validation. Furthermore, we introduce the Knowledge-Guided STL Transformation (KGST) framework, a novel approach for transforming natural language into STL, involving a generate-then-refine process based on external knowledge. Statistical analysis shows that the STL-DivEn dataset exhibits more diversity than the existing NL-STL dataset. Moreover, both metric-based and human evaluations indicate that our KGST approach outperforms baseline models in transformation accuracy on STL-DivEn and DeepSTL datasets.  [^1] Project can be found at https://github.com/YueFang0618/STL-DivEn
@@
Yue Fang 1, Zhi Jin 1, Jie An 2, Hongshen Chen 3, Xiaohong Chen 4, and Naijun Zhan 1 1Peking University, Beijing, China 2Institute of Software, Chinese Academy of Sciences, Beijing, China 3JD.com, Beijing, China 4East China Normal University, Shanghai, China y.fang@stu.pku.edu.com,zhijin@pku.edu.cn
@@ 1 Introduction
Writing accurate STL formulas directly is a huge burden for domain experts and even for researchers, as it is both time-consuming and error-prone. With the development of natural language processing (NLP) technology, researchers have been experimenting with utilizing NLP technology to transform natural language into TL and STL expressions, aiming to improve its automation and accuracy. For example, Lignos et al. (2015); Ghosh et al. (2016) use predefined pattern formulas to transform natural language sentences into an intermediate representation. Subsequently, by applying a set of predefined rules manually, the intermediate representation is mapped to temporal logic formulas. These approaches require extensive domain expertise and involve a steep learning curve Kulkarni et al. (2013). Specifically, they can only be applied to very restrictive structured natural language expressions that match the given patterns.
@@
  • The proposed KGST framework demonstrates superior performance not only on the newly developed STL-DivEn dataset but also on the existing DeepSTL dataset. This highlights its versatility and robustness across different datasets.
@@ 6 Conclusion
To address this issue, at least the following approaches can be considered in the future. First, we can extract temporal property patterns from existing real-world cyber-physical systems. Second, for specific domains like autonomous driving, we can extract necessary data from domain-related requirements documentation, e.g., international standards related to AUTOSAR for electronic vehicles. Furthermore, we can infer possible timing properties and other temporal characteristics of cyber-physical systems by simulating their real interactive environments. In this way, our dataset can be continuously enriched by incorporating human validation to train better models.

### v2

@@ Abstract
Temporal Logic (TL), especially Signal Temporal Logic (STL), enables precise formal specification, making it widely used in cyber-physical systems such as autonomous driving and robotics. Automatically transforming NL into STL is an attractive approach to overcome the limitations of manual transformation, which is time-consuming and error-prone. However, due to the lack of datasets, automatic transformation currently faces significant challenges and has not been fully explored. In this paper, we propose a NL-STL dataset named STL-Diversity-Enhanced (STL-DivEn), comprising 16,000 samples enriched with diverse patterns. To develop the dataset, we first manually create a small-scale seed set of NL-STL pairs. Next, representative examples are identified through clustering and used to guide large language models (LLMs) in generating additional NL-STL pairs. Finally, diversity and accuracy are ensured through rigorous rule-based filters and human validation. Furthermore, we introduce the Knowledge-Guided STL Transformation (KGST) framework, a novel approach for transforming natural language into STL, involving a generate-then-refine process based on external knowledge. Statistical analysis shows that the STL-DivEn dataset exhibits more diversity than the existing NL-STL dataset. Moreover, both metric-based and human evaluations indicate that our KGST approach outperforms baseline models in transformation accuracy on STL-DivEn and DeepSTL datasets.  [^1] Project can be found at https://github.com/YueFang0618/STL-DivEn
@@
Yue Fang1,2, Zhi Jin1,2thanks: Corresponding author, Jie An311footnotemark: 1, Hongshen Chen4, Xiaohong Chen5, Naijun Zhan1,2 1School of Computer Science, Peking University, Beijing, China 2Key Laboratory of High Confidence Software Technologies (PKU), MOE, China 3National Key Laboratory of Space Integrated Information System, Institute of Software, Chinese Academy of Sciences, Beijing, China 4JD.com, Beijing, China 5East China Normal University, Shanghai, China y.fang@stu.pku.edu.cn, zhijin@pku.edu.cn, anjie@iscas.ac.cn
@@ 1 Introduction
Writing accurate STL formulas directly is a huge burden for domain experts, as it is both time-consuming and error-prone. With the development of natural language processing (NLP) technology, researchers have been experimenting with utilizing NLP technology to transform natural language into TL and STL expressions, aiming to improve its automation and accuracy. For example, Lignos et al. (2015); Ghosh et al. (2016) use predefined pattern formulas to transform natural language sentences into an intermediate representation. Subsequently, by applying a set of predefined rules manually, the intermediate representation is mapped to temporal logic formulas. These approaches require extensive domain expertise and involve a steep learning curve Kulkarni et al. (2013). Specifically, they can only be applied to very restrictive structured natural language expressions that match the given patterns.
@@
  • The proposed KGST framework demonstrates superior performance not only on the STL-DivEn dataset but also on the existing DeepSTL dataset. This highlights its versatility and robustness across different datasets.
@@ 6 Conclusion
To address this issue, at least the following approaches can be considered in the future. First, we can extract temporal property patterns from existing real-world cyber-physical systems. Second, for specific domains like autonomous driving, we can extract the necessary data from domain-related requirements documentation, e.g., international standards related to AUTOSAR for electronic vehicles. Furthermore, we can infer possible timing properties and other temporal characteristics of cyber-physical systems by simulating their real interactive environments. In this way, our dataset can be continuously enriched by incorporating human validation to train better models.
Acknowledgement
We sincerely thank the anonymous reviewers for their valuable comments and suggestions. This work is supported by the National Natural Science Foundation of China under Grant Nos. 62192731, 62192732, and 62192730.